OpenStudy (anonymous):

How do you make a table of values with the equation y=3x-2 with the replacement set {-2, -1, 0, 1, 2}? (Please respond quickly. I have no idea how to solve this.)

jimthompson5910 (jim_thompson5910):

For each number in the set, you'll replace x with that value example: the first number in the set is -2. Replace x with -2 and evaluate to find y y = 3x - 2 y = 3*(-2) - 2 ... replace x with -2 y = -6-2 y = -8 So you'll have one row in your x,y table that has -2 under the x column and -8 under the y column